OpenStudy (anonymous):

The sum of three consecutive even integers is -24. Find the integers.

OpenStudy (watchmath):

The same idea spoyled. Why don't you give it a try first? :D

OpenStudy (anonymous):

I dont understand this at all. Hoping to look at these and try to comprehend. These are all my practice exercises.

OpenStudy (watchmath):

Ok, if the smallest number among the three is \(n\). Then the next even number is \(n+2\) and the next after that is \(n+2+2=n+4\). So the sum of these three even numbers is \(n+(n+2)+(n+4)=3n+6\) we want this sum to be equal to -24. So \(3n+6=-24\) \(3n=-30\) \(n=-10\) So the next two even numbers are -8, -6. Thus -10,-8,-6 are the answers.

OpenStudy (radar):

Les 2x = 1st integer 2x + 2 = 2nd 2x+4 = 3rd even integer the sum is -24 so 2x+2x+2+2x+4=-24 6x=-24-6=-30 x=-5 2x=-10 2x+2=-8 2x+4=-6 Making first digit equal 2x assures it is even. the next one is also even so it is 2x+2

OpenStudy (anonymous):

thank you and that helps to understand process little better